Please help! The caterers for a semi formal dinner and dance are preparing two different meals; chicken at 12$ or pasta at 8$. The total cost of the dinners for 240 students is $2100.
A. Create a linear system to model this situation. Define your variables.
B. Solve the system to determine the number of chicken and pasta dinners the students ordered

Answer by macston(5194)

.
A. Create a linear system to model this situation. Define your variables.
.
C=number of chicken plates; P=number of pasta plates
.
C+P=240
$12C+$8P=$2100
.
B. Solve the system to determine the number of chicken and pasta dinners the students ordered
.
C+P=240
P=240-C . (Use this to substitute for P below)
.
$12C+$8P=$2100 . (Substitute for P - see above)
$12C+$8(240-C)=$2100
$12C+$1920-$8C=$2100
$4C=$180
C=45
ANSWER 1: They ordered 45 chicken dinners.
.
P=240-C=240-45=195
ANSWER 2: They ordered 195 pasta dinners.
.
CHECK:
$12C+$8P=$2100
$12(45)+$8(195)=$2100
$540+$1560=$2100
$2100=$2100
.